OpenStudy (anonymous):

What is the sum of the geometric sequence 3, 15, 75, … if there are 7 terms?

OpenStudy (anonymous):

use sum=a1(r^(n-1)-1)/(r-1)

OpenStudy (anonymous):

how do i plug it in

OpenStudy (anonymous):

sorry it shud be r^n instead of R^9n-1) n= no. of terms r= common ratio a1=first term

OpenStudy (anonymous):

thkss

OpenStudy (anonymous):

answer shud be 58593

OpenStudy (anonymous):

S7 = 3 * 5^7 - 1 ------ 5 - 1 = 58593